# Machine Learning Nonadiabatic Dynamics: Eliminating Phase Freedom of Nonadiabatic Couplings with the State-Interaction State-Averaged Spin-Restricted Ensemble-Referenced Kohn–Sham Approach

### v1

@@ 3.1 Model evaluation
We first evaluated the models for their ability to predict diabatic properties, specifically $E^{PPS}$, $E^{OSS}$, and $\Delta^{2}$. A database was generated consisting of 48,750 data points, randomly selected from 50 surface-hopping trajectories obtained using exact factorization-based dynamics (SHXF)21. These trajectories were computed using the SSR(2,2) formalism combined with the $\omega$PBEh67/6-31g*68 level of theory, with a time step of 0.24 fs over a total simulation duration of 300 fs.
@@
$E^{PPS}$ 0.041 0.074 0.999 0.065 0.245 0.999
$E^{OSS}$ 0.057 0.102 0.999 0.081 0.314 0.999
$\Delta^{2}$ 0.564 0.978 0.998 1.370 5.840 0.992
Table 1: Mean Absolute Error (MAE), Root-Mean-Square Error (RMSE), and R2 values of the trained NequIP-BAM model. $E^{PPS}$ and $E^{OSS}$ are given in kcal/mol, and $\Delta^{2}$ values are in (kcal/mol)2.
@@
As shown in Table 1 and Figure 2, the MLP predictions for diabatic properties are highly accurate. The MAE for the energy predictions of $E^{PPS}$ and $E^{OSS}$ is approximately 0.041 and 0.057 kcal/mol, respectively, while the gradient errors are about 0.065 and 0.081 kcal/mol/Å. Both $E^{PPS}$ and $E^{OSS}$ show high accuracy, surpassing the threshold for chemical accuracy. However, the error for $E^{OSS}$ is slightly larger than that of $E^{PPS}$, likely due to the greater difficulty in mapping excited-state PES from molecular geometry compared to ground-state PES.
@@
Increasing the number of layers in the Interaction Block could potentially enhance accuracy. Indeed, when we tested the model’s accuracy with varying numbers of layers, we observed nearly a twofold improvement in accuracy with the addition of an extra layer for $E^{OSS}$. We also compared the model with different numbers of feature channels, and the results showed that increasing the number of features was less effective compared to increasing the number of layers. However, due to computational cost constraints, we used 5 layers and 64 feature channels in our final model. Supplementary Table S1 presents the results for different layer configurations, while Supplementary Table S2 provides the results for various feature channels. For $\Delta^{2}$, the MAE of the energy is approximately 0.564 (kcal/mol)2 for the NequIP-BAM model. Our approach is notable not only for its accuracy but also for successfully resolving the scattered pattern observed in previous work58, which used a phaseless loss function. This result demonstrates that squaring the double-valued $\Delta$ is an effective strategy for generating a continuous, smooth function that can be efficiently trained. This suggests that the model is no longer affected by the double-valued phase problem and can predict nonadiabatic properties with a similar level of accuracy to conventional models.
@@
Next, we compared the effects of different activation functions applied to the bias layer. During $\Delta^{2}$ training, we observed negative output values from MLP, which are undesirable given the positive nature of the $\Delta^{2}$. Therefore, ReLU and SiLU activation functions were applied during the $\Delta^{2}$ training and these results are presented in Table 2. The ReLU activation function improves the accuracy compared against the model without the activation function. The MAE of $\Delta^{2}$ with ReLU activation is 0.564 (kcal/mol)2, compared to 0.646 (kcal/mol)2 for the model with the identity activation. When the reference value was near zero value, using the identity activation function occasionally resulted in negative predictions for $\Delta^{2}$, which are theoretically impossible. In contrast, the ReLU activation function produced positive predictions due to its bias. This suggests that ReLU is effective in reducing prediction errors for $\Delta^{2}$. While ReLU showed a smaller test error compared to the identity function, the SiLU activation resulted in a significantly larger error of 26.686 (kcal/mol)2. This is likely due to SiLU’s handling of negative values, designed to address the dying ReLU problem, which appears incompatible with the strict positive constraint of $\Delta^{2}$. Based on these findings, we concluded that the ReLU activation function, combined with the NequIP-BAM model, is the most suitable MLP architecture for training PSB3 model. This architecture was applied in all subsequent training and ML-ESMD simulations.
@@
Identity 0.646 1.239 0.997 1.100 4.785 0.994
ReLU 0.564 0.978 0.998 1.370 5.840 0.992
SiLU 26.686 33.935 0.080 2.568 5.411 0.991
@@
Table 2: Mean Absolute Error (MAE), Root-Mean-Square Error (RMSE), and R2 values of the trained $\Delta^{2}$ model. The units are expressed in (kcal/mol)2.
@@
While overall model performance is important, accurately predicting the minimum-energy conical intersection (MECI) structure is critical due to its central role in nonadiabatic dynamics involving multiple electronic states. The degeneracy of the two adiabatic surfaces at the CI is lifted by two branching plane vectors: the difference gradient vector $\boldsymbol{g}$ and the coupling derivative vector $\boldsymbol{h}$. To assess the model’s accuracy at the CI, we compared the predicted branching plane vectors and MECI structures with their respective references. Figure 3 shows the MECI structures and two branching plane vectors for the CI structure, as predicted by the NequIP-BAM model and calculated by SSR(2,2)/$\omega$PBEh/6-31G*. As shown in Figure 3(a), the MECI structures predicted by the NequIP-BAM model and calculated using SSR(2,2) are nearly indistinguishable. The $\boldsymbol{g}$-vector represents the bond elongation motion of the terminal C=C bond, while the $\boldsymbol{h}$-vector corresponds to the torsional motion of the central C=C bond. These two motions are known to significantly contribute to molecular dynamics at the CI in PSB3, as previously reported69. The predicted vectors at the CI region match the direction and magnitude of the reference values.
@@
Figure 3: MECI structure and its branching plane vectors: (a) Comparison of MECI structures predicted by the NequIP-BAM model (blue) and calculated by SSR(2,2)/$\omega$PBEh/6-31G* (red). (b) difference gradient vector ($\vec{g}$) and (a) coupling derivative vector ($\vec{h}$). The red and blue arrows denote vectors calculated by SSR(2,2)/$\omega$PBEh/6-31G and predicted by the NequIP-BAM model respectively.
@@
Since NACV plays a key role in nonadiabatic dynamics, we evaluated the prediction accuracy of NACV in the vicinity of the CI region by comparing the NequIP-BAM model with the reference SSR(2,2) values. NACVs were calculated using eq 8, with all components computed from the MLP predictions of $E^{PPS}$, $E^{OSS}$, $\Delta^{2}$, and their respective gradients. We divided the NACV test set into two groups: one near the CI region and one away from it (Figure 4). Each group of molecules is composed of 404 and 2,285 molecules, respectively, each containing 42 elements. Thus, Figure 4 includes 16968 elements in the vicinity of the CI and 95970 elements distant from the CI. The division was based on an adiabatic energy difference of 0.5 eV. Both positive and negative vector elements were predicted, matching their reference values, unaffected by the double-valued problem. Accurate NACV prediction near the CI is crucial for describing correct state transitions. Additionally, accurate prediction of NACVs far from the CI region is also important to prevent erroneous transitions in regions where the coupling between electronic states is weak. We confirm that the NACV predictions are not overestimated and remain accurate when it’s away from the CI (Figure 4). In surface hopping dynamics, where hopping probabilities are directly influenced by NACVs, the precise prediction of NACVs by our MLP strengthens the reliability of state transitions in our ML-ESMD simulations.
@@ 3.2 Nonadiabatic molecular dynamics
After successfully demonstrating that our MLP model accurately reproduces static properties, we ran nonadiabatic dynamics simulations. We used MLP for nonadiabatic MD simulations based on the SHXF method21, implemented in the PyUNIxMD package70, 71. For the initial conditions, 100 nuclear configurations and their velocities were sampled from the optimized ground-state geometry of the $trans$-PSB3 molecule, randomly selected from Wigner distribution on 300K. Dynamics were performed over a total of 300 fs with a timestep of 0.24 fs. The nonadiabatic relaxation of PSB3 from the $S_{1}$ state to the ground state via MECI involves changes in the torsion angle of the central C=C bond, which leads to the differentiation to the $cis$ and $trans$ isomers69.
We tracked the dihedral angle of the central C=C bond (Figure 5). The dihedral angle trajectories clearly show the separation into the $cis$ and $trans$ isomers. Although the results from 100 trajectories may not be statistically significant, the final cis-to-trans ratio of 61:39 is close to the reference ratio of 63:3769. Current work improves on the previous work58, which reported a ratio of 58.5:41.5. Additionally, the comparison of ML-based and SSR-based trajectories reveals that the molecular motions and final yields are in strong agreement between the two approaches (Figure 4 and 5).
We also confirm that the average electronic population evolutions from ML and reference SSR dynamics simulations are in agreement (Figure 5(b)). Averaged populations $P_{i}$ and $p_{i}$ are calculated from the running state and BO population, where $P_{i}=N_{i}/N_{traj}$ and $p_{i}=\sum\limits_{I}^{N_{traj}}\rho_{ii}(t)/N_{traj}$, respectively. Here, $N_{i}$, $N_{traj}$, and $\rho_{ii}$ represent the number of trajectories running on the $i$-th electronic state, the total number of trajectories, and a BO population at time $t$, respectively. The agreement between the two populations highlights the MLP’s ability to accurately capture electron-nuclear correlations. Furthermore, the population trends observed in our SHXF/ML dynamics closely match those from the SHXF/SSR(2,2) simulations, despite the ML model being approximately 100 times faster.
@@ 4 Conclusion
We propose a phaseless ML-ESMD approach by building MLPs for the diagonal Hamiltonian elements—$E^{PPS}$, $E^{OSS}$, and $\Delta^{2}$—obtained from the SSR (2,2) method.
This formalism addresses the phase arbitrariness in ML-ESMD. By focusing on predicting diabatic Hamiltonian elements rather than directly on the NACV, we avoid the singularities at CIs. Additionally, we reformulate the NACV and its gradients by replacing the phase-dependent off-diagonal elements with their squared values $\Delta^{2}$. MLP models trained with this approach show high accuracy without any discontinuities or divergence. We show that the ML-ESMD of the PSB3 system agrees well with the reference ab initio simulations. We expect that this approach, when combined with state-of-the-art MLPs, will aid in the development of universal excited-state MLPs. These models could be integrated into de novo molecular design, allowing for the exploration of ever more complex and challenging problems in the field.
@@ 5 Code availability
This research was supported by the National Research Foundation of Korea (NRF) funded by the Korean government (Ministry of Science and ICT(MSIT))(NRF-2023M3K5A1094813, RS-2023-00257666, RS-2024-00455131, NRF-2022R1C1C1010605) and by Institute for Basic Science (IBS-R036-D1). SYW and CWM acknowledge the support from Brain Pool program funded by the Ministry of Science and ICT through the National Research Foundation of Korea (No. RS-2024-00407680). CWM, THP, and SYW are grateful for the computational support from the Korea Institute of Science and Technology Information (KISTI) for the Nurion cluster (KSC-2023-CRE-0502, KSC-2023-CRE-0355, KSC-2023-CRE-0059, KSC-2022-CRE-0424, KSC-2022-CRE-0115, KSC-2022-CRE-0113, KSC-2021-CRE-0542, KSC-2021-CRE-0496, KSC-2022-CRE-0429, KSC-2022-CRE-0469, KSC-2023-CRE-0050, KSC-2023-CRE-0251, KSC-2023-CRE-0261, KSC-2023-CRE-0332, KSC-2023-CRE-0472, KSC-2023-CRE-0501, KSC-2024-CRE-0358). {suppinfo} See the supplementary material for a detailed compilation of the obtained results as well as further data and analysis to support the points made throughout the text.

### v2

@@ 3.1 Model evaluation
We first evaluated the models for their ability to predict diabatic properties, specifically $E^{PPS}$, $E^{OSS}$, and $\Delta^{2}$. As shown in Table 1 and Figure 2, the MLP predictions for diabatic properties are highly accurate. The MAE for the energy predictions of $E^{PPS}$ and $E^{OSS}$ is approximately 0.040 and 0.058 kcal/mol, respectively, while the gradient errors are about 0.063 and 0.080 kcal/mol/Å. Both $E^{PPS}$ and $E^{OSS}$ show high accuracy, surpassing the threshold for chemical accuracy. However, the error for $E^{OSS}$ is slightly larger than that of $E^{PPS}$, as shown in Table 1 and Figure S1, likely due to the greater difficulty in mapping excited-state PES from molecular geometry compared to ground-state PES.
@@
$E^{PPS}$ 0.040 0.072 0.999 0.063 0.181 0.999
$E^{OSS}$ 0.058 0.111 0.999 0.080 0.259 0.999
$\Delta^{2}$ 0.463 0.810 0.999 1.135 2.626 0.998
Table 1: Mean Absolute Error (MAE), Root-Mean-Square Error (RMSE), and R2 values of the NequIP-BAM on the test dataset. $E^{PPS}$ and $E^{OSS}$ are given in kcal/mol, and $\Delta^{2}$ values are in (kcal/mol)2.
@@
Increasing the number of layers in the Interaction Block could potentially enhance accuracy. Indeed, when we tested the model’s accuracy with varying numbers of layers, we observed nearly a twofold improvement in accuracy with the addition of an extra layer for $E^{OSS}$. We also compared the model with different numbers of feature channels, and the results showed that increasing the number of features was less effective compared to increasing the number of layers. However, due to computational cost constraints, we used 5 layers and 64 feature channels in our final model. Table S1 presents the results for different layer configurations, while Table S2 provides the results for various feature channels. For $\Delta^{2}$, the MAE of the energy is approximately 0.463 (kcal/mol)2 for the NequIP-BAM model. In addition, we observe a larger error for the small $\Delta^{2}$ region (Figure S1(c)). This occurs because $\Delta^{2}$ becomes small near CIs as well as in the region far away from CIs with a large energy gap $|E_{PPS}-E_{OSS}|$. Thus, vastly different molecular geometries can result in similar $\Delta^{2}\approx 0$ values, making it inherently more challenging for the MLP to distinguish between these geometries. Our approach is notable not only for its accuracy but also for successfully resolving the scattered pattern observed in previous work62, which used a phaseless loss function. This result demonstrates that squaring the double-valued $\Delta$ is an effective strategy for generating a continuous, smooth function that can be efficiently trained. This suggests that the model is no longer affected by the double-valued phase problem and can predict nonadiabatic properties with a similar level of accuracy to conventional models.
@@
Next, we compared the effects of different activation functions applied to the bias layer. During $\Delta^{2}$ training, we observed negative output values from MLP, which are undesirable given the positive nature of the $\Delta^{2}$. Therefore, ReLU and SiLU activation functions were applied during the $\Delta^{2}$ training and these results are presented in Table 2. The ReLU activation function improves the accuracy compared against the model without the activation function. The MAE of $\Delta^{2}$ with ReLU activation is 0.463 (kcal/mol)2, compared to 0.634 (kcal/mol)2 for the model with the identity activation. When the reference value was near zero value, using the identity activation function occasionally resulted in negative predictions for $\Delta^{2}$, which are theoretically impossible. In contrast, the ReLU activation function produced positive predictions due to its bias. This suggests that ReLU is effective in reducing prediction errors for $\Delta^{2}$. While ReLU showed a smaller test error compared to the identity function, the SiLU activation resulted in a significantly larger error of 29.090 (kcal/mol)2. The issue likely comes from SiLU’s handling of negative values. It is designed to prevent the dying ReLU problem, where neurons become inactive and gradients vanish78. However, this conflicts with the strict requirement for $\Delta^{2}$ to remain positive. Based on these findings, we concluded that the ReLU activation function, combined with the NequIP-BAM model, is the most suitable MLP architecture for training PSB3 model. This architecture was applied in all subsequent training and ML-ESMD simulations.
@@
Identity 0.634 1.084 0.998 1.124 2.186 0.999
ReLU 0.463 0.810 0.999 1.135 2.626 0.998
SiLU 29.090 36.472 0.116 2.723 4.961 0.994
@@
Table 2: Mean Absolute Error (MAE), Root-Mean-Square Error (RMSE), and R2 values of the NequIP-BAM for $\Delta^{2}$ prediction on the test dataset. The units are expressed in (kcal/mol)2.
@@
While overall model performance is important, accurately predicting the minimum-energy conical intersection (MECI) structure is critical due to its central role in nonadiabatic dynamics involving multiple electronic states. The degeneracy of the two adiabatic surfaces at the CI is lifted by two branching plane vectors: the difference gradient vector $\vec{g}$ and the coupling derivative vector $\vec{h}$. To assess the model’s accuracy at the CI, we compared the predicted branching plane vectors and MECI structures with their respective references. Figure 3 shows the MECI structures and two branching plane vectors for the CI structure, as predicted by the NequIP-BAM model and calculated by SSR(2,2)/$\omega$PBEh/6-31G* (See Supporting Information S9 for the geometry of CI and two branching plane vectors). As shown in Figure 3(a), the MECI structures predicted by the NequIP-BAM model and calculated using SSR(2,2) are nearly indistinguishable. The $\vec{g}$-vector represents the bond elongation motion of the terminal C=C bond, while the $\vec{h}$-vector corresponds to the torsional motion of the central C=C bond. These two motions are known to significantly contribute to molecular dynamics at the CI in PSB3, as previously reported68. The predicted vectors at the CI region match the direction and magnitude of the reference values. In addition, we can also predict the adiabatic potential energy surface and discontinuity of real-valued eigenvectors around the CI in the branching plane (Figure S2), which is critical to assess Berry phase due to conical topology of CIs.
@@
Figure 3: MECI structure and its branching plane vectors: (a) Comparison of MECI structures predicted by the NequIP-BAM model (blue) and calculated by SSR(2,2)/$\omega$PBEh/6-31G* (red). (b) difference gradient vector ($\vec{g}$) and (a) coupling derivative vector ($\vec{h}$). The red and blue arrows denote vectors calculated by SSR(2,2)/$\omega$PBEh/6-31G* and predicted by the NequIP-BAM model respectively.
@@
Since NACV plays a key role in nonadiabatic dynamics, we evaluated the prediction accuracy of NACV in the vicinity of the CI region by comparing the NequIP-BAM model with the reference SSR(2,2) values. NACVs were calculated using eq 8, with all components computed from the MLP predictions of $E^{PPS}$, $E^{OSS}$, $\Delta^{2}$, and their respective gradients. We divided the NACV test set into two groups: one near the CI region (404 geometries) and one away from it (2,285 geometries)(Figure 4). The NACV vector contains 42 components. Thus, Figure 4 includes 16968 elements in the vicinity of the CI and 95970 elements distant from the CI. The division was based on an adiabatic energy difference of 0.5 eV. The predicted vector elements from the model closely matched their reference values, with small errors for most points (Figure S1(d)). Accurate NACV prediction near the CI is crucial for describing correct state transitions. Additionally, accurate prediction of NACVs far from the CI region is also important to prevent erroneous transitions in regions where the coupling between electronic states is weak. We confirm that the NACV predictions are not overestimated and remain accurate when it’s away from the CI (Figure 4). In surface hopping dynamics, where hopping probabilities are directly influenced by NACVs, the precise prediction of NACVs by our MLP strengthens the reliability of state transitions in our ML-ESMD simulations.
@@ 3.2 Nonadiabatic molecular dynamics
After successfully demonstrating that our MLP model accurately reproduces static properties, we ran nonadiabatic dynamics simulations. We used MLP for nonadiabatic MD simulations based on the SHXF method21, implemented in the PyUNIxMD package79, 80. For the initial conditions, 100 nuclear configurations and their velocities were newly sampled from the optimized ground-state geometry of the $trans$-PSB3 molecule, randomly selected from Wigner distribution on 300K. Dynamics were performed over a total of 300 fs with a timestep of 0.24 fs. The nonadiabatic relaxation of PSB3 from the $S_{1}$ state to the ground state via MECI involves changes in the torsion angle of the central C=C bond, which leads to the differentiation to the $cis$ and $trans$ isomers68.
We tracked the dihedral angle of the central C=C bond (Figure 5). The dihedral angle trajectories clearly show the separation into the $cis$ and $trans$ isomers. Although the results from 100 trajectories may not be statistically significant, the final cis-to-trans ratio of 61:39 is close to the reference ratio of 63:3768. Current work improves on the previous work62, which reported a ratio of 58.5:41.5. Additionally, the comparison of ML-based and SSR-based trajectories reveals that the molecular motions and final yields are in strong agreement between the two approaches (Figure 4 and 5).
We also confirm that the average electronic population evolutions from ML and reference SSR dynamics simulations are in agreement (Figure 5(b)). As described in Figure S3, the magnitude of nonadiabatic couplings from ML prediction increases properly as the two adiabatic energies close to each other. Averaged populations $P_{i}$ and $p_{i}$ are calculated from the running state and BO population, where $P_{i}=N_{i}/N_{traj}$ and $p_{i}=\sum\limits_{I}^{N_{traj}}\rho_{ii}(t)/N_{traj}$, respectively. Here, $N_{i}$, $N_{traj}$, and $\rho_{ii}$ represent the number of trajectories running on the $i$-th electronic state, the total number of trajectories, and a BO population at time $t$, respectively. The agreement between the two populations highlights the MLP’s ability to accurately capture electron-nuclear correlations. Furthermore, the population trends observed in our SHXF/ML dynamics closely match those from the SHXF/SSR(2,2) simulations, despite the ML model being approximately 100 times faster.
@@ 4 Conclusion
We propose a phaseless ML-ESMD approach by building MLPs for the diabatic Hamiltonian elements—$E^{PPS}$, $E^{OSS}$, and $\Delta^{2}$—obtained from the SSR (2,2) method. This formalism addresses the phase arbitrariness in ML-ESMD. By focusing on predicting diabatic Hamiltonian elements rather than directly on the NACV, we avoid the singularities at CIs. Additionally, we reformulate the NACV and its gradients by replacing the phase-dependent off-diagonal elements with their squared values $\Delta^{2}$. MLP models trained with this approach show high accuracy without any discontinuities or divergence. We show that the ML-ESMD of the PSB3 system agrees well with the reference ab initio simulations. We expect that our ML approach, when combined with state-of-the-art MLPs, will aid in the development of multi-state universal MLPs, applicable to systems encompassing elements across the entire periodic table, which remains as future work.81 These models could be integrated into de novo molecular design, allowing for the exploration of ever more complex and challenging problems in the field.
@@ 5 Code availability
This research was supported by the National Research Foundation of Korea (NRF) funded by the Korean government (Ministry of Science and ICT(MSIT))(NRF-2023M3K5A1094813, RS-2023-00257666, RS-2024-00455131, NRF-2022R1C1C1010605) and by Institute for Basic Science (IBS-R036-D1). SYW and CWM acknowledge the support from Brain Pool program funded by the Ministry of Science and ICT through the National Research Foundation of Korea (No. RS-2024-00407680). CWM, THP, and SYW are grateful for the computational support from the Korea Institute of Science and Technology Information (KISTI) for the Nurion cluster (KSC-2023-CRE-0502, KSC-2023-CRE-0355, KSC-2023-CRE-0059, KSC-2022-CRE-0424, KSC-2022-CRE-0115, KSC-2022-CRE-0113, KSC-2021-CRE-0542, KSC-2021-CRE-0496, KSC-2022-CRE-0429, KSC-2022-CRE-0469, KSC-2023-CRE-0050, KSC-2023-CRE-0251, KSC-2023-CRE-0261, KSC-2023-CRE-0332, KSC-2023-CRE-0472, KSC-2023-CRE-0501, KSC-2024-CRE-0358). Computational work for this research was partially performed on the Olaf supercomputer supported by IBS Research Solution Center. {suppinfo} See the supplementary material for a detailed compilation of the obtained results as well as further data and analysis to support the points made throughout the text.